shazam288:

What is the product? (6 r minus 1)(negative 8 r minus 3)

MrMudd183:

what's the first thing you think should be done?

MrMudd183:

\[6r-1=-8r-3\] is this how the problem is set up? if so then the first thing i would do is add the -1 to the -3 then the problem would be \[6r=-8r-2\] after that you would add -8r to 6r to end up with \[14r=-2\] finally you would divide both sides by 14 and the answer should be r=-0.143

MrMudd183:

That's what I got

Mercury:

The two quantities are being multiplied, not set equal to each other. Hence, the word "product." To evaluate this, you need to apply FOIL (First, outer, inner, last) (a+b)(c+d) = a*c + a*d + b*c + b*d notice how the first term in each bracket (a and c) get multiplied first, then the "outer" terms (a and d), etc. and all the products are added up. can you try applying this to your problem? be careful with signs.

jhonyy9:

exactly @Mercury ,i dont know @MrMudd183 how you get this sign of equality between these 2 expressions inside parentheses ?

MrMudd183:

I thought of it wrong, my bad

MrMudd183:

i have geomerty this semester and my mind went straight to that, sorry

jhonyy9:

@Mercury please there inside these parentheses where you wrote ,,outer" dont need being b and d bc. you wrote a and d ???

Mercury:

... what? in FOIL, the outer terms *would* be a and d  https://encrypted-tbn0.gstatic.com/images?q=tbn%3AANd9GcSs9_oROzDef2KN9zmTN5_BOwEp_fXWUh4nCwJfgVZMzA&usqp=CAU&ec=45690269

Mercury:

in this case, "outer" doesn't mean "outside the parenthesis" it means the outermost terms from the center